OpenStudy (anonymous):

Sam's bike tire contains 15 units of air particles and has a volume of 160mL. Under these conditions the pressure reads 13 psi. The tire develops a leak. Now it contains 10 units of air particles and has contracted to a volume of 150mL. What would the tire pressure be now?

OpenStudy (aaronq):

PV=nRT one system equals another: PV/nRT=PV/NRT remove what is being kept constant: \[\frac{P _{1}V _{1}}{ n _{1} } = \frac{P _{2}V _{2}}{ n _{2} } \]